# DropleX: Liquid sensing on tablet touchscreens

### v3

@@ 8.2. Concentration detection
We find that in the case of DNA and salt water (Fig. 17a,c) our system is primarily effective as a binary detector of trace concentrations of liquid. Specifically in the case of DNA it is able to consider concentration of 0–20 ng/ as a low concentration and 40 ng/ as a high concentration and distinguish between those, essentially functioning as a detector with a threshold of 20 ng/, yielding an overall detection accuracy of 95.6%. Similarly, in the case of salt water, the touchscreen is able to function a detector of the liquid with a threshold of M with an overall detection accuracy of 95.7%. In the case of ethanol, we find that our system is able to distinguish between six different concentrations of alcohol specifically 0, 10, 30, 50, 80, 100% with an overall accuracy of 93% (Fig. 17b).
@@ 8.4. Benchmarks
Measurement stability over time. We evaluate the stability of capacitance readings for the priming drop and the test liquid samples on the tablet screen. We show in Fig. 20a that the measured capacitance at the centroid of the priming drop remains stable across the measured period of ten minutes with a mean and standard deviation of 376 7 device units. We also perform this measurement for water samples of different volumes across an 10 minute period and show in Fig. 20b that these values remain stable across the measured period with standard deviations ranging from 6 to 10 device units across the measurement period. There is no substantial relationship between the standard deviation of the measurements with the volume of the liquid drops. We note that stability across these time scales is sufficient for averaging the effects of noise and for recording a measurement.
Effect of temperature. We examine the effect of temperature on measured capacitance. We perform this experiment using tap water and deionized water. We heated up the liquid samples in a tabletop water bath (Fig. 21a) (ivy, 2025) from 30–90∘C. The results in Fig. 21b show that the measured capacitance increases for both liquids. We note that the reason is different for both liquids tested. In the case of tap water which has more ions than deionized water, the ions mobility will increase which will increase the conductivity, and thus a larger capacitance. In contrast, with deionized water, where there are no ions, the capacitance increase is due to changes in the permittivity which increases the capacitance. We note that in this experiment, the measured capacitance for the tap water increases to approximately -800 device units, which is the empirical limit of the system.
@@
Effect of device tilt. As the user may not always have their tablet flat on a surface, particularly in field conditions outdoors, the effect of uneven surfaces on our system. To do this, we placed the tablet on a motorized tilt platform (til, 2025) (Fig. 23a) and adhered the tablet to it using some putty. We deposited ten drops of water and isopropyl alcohol (IPA) onto the surface and varied the tilt angle at increments of approximately 1 to 3∘. We show in Fig. 23b how the droplets displace from their original position as tilt angle increases. In the case of water, which has a surface tension of , the droplets start moving at 5∘ and this displacement increases till its roll-off angle of 11.5∘ when it slides off the tablet. In contrast, IPA, which has a lower surface tension of , has a much higher rate of displacement with a roll-off angle of 2∘. While water had a displacement of approximately 4 cells at the roll-off angle, alcohol had a displacement around three times higher at 12 cells. The behavior of the drops at the rest and roll-off angles are visualized in Fig. 23c.
Ablation study. We perform an ablation study on the liquid adulteration datasets to examine the effect of different system design choices on performance (Fig. 24). Specifically, we examine the effect of number of training samples, region size around the centroid, and machine learning model on system performance. For these experiments, we consider a subset of the adulteration classes: for soda spiking and wine adulteration we compared to ethanol of 10, 20, 30% as a matter of computational efficiency to examine the overall effect of the system parameters. Here we used 500 drop sizes.
For the training dataset size (Fig. 24a) we evaluated the affect of changing the size of the training set from 0, 25, 50, 75, 100% of the original size. We find that as expected this increases the accuracies with diminishing returns. In the soda and wine dataset, we find that there are no substantial performance increases beyond approximately 200 training samples.
For the region size (Fig. 24b), we find that accuracy increases steadily as the region size increases from a patch to a patch. In the case we use a fully-connected network instead of a CNN. For larger classes we use an adaptive CNN where for and classes there are only two convolutional layers without any pooling, while for larger patch sizes of to there is a single max pooling layer, and for there are 2 pooling layers. This result is expected as the screen area covered by a drop can expand to affect the electric field of a grid, and a larger grid captures this information.
Finally, we examined the effect of model architecture on system performance (Fig. 24c). We find that using a CNN only, LSTM only, CNN+LSTM and CNN+Transformer provide performance within a similar range of 91 to 94%. Tree-based models such as XGBoost and Random Forest provide lowered accuracies of 79 to 81%. The results suggest that global features about the region, can be captured by the tree-based models. However, deep learning modules better capture the non-linear relationship (Sec. 5) between classes, with the CNN and LSTM producing higher accuracies. As CNN alone and LSTM alone are able to produce high accuracies, it suggests that non-linear feature combinations are sufficient to perform classification.
@@ 9.1. Limitations
Geometry and thickness of container base. The containers that we selected for our through-container sensing experiments had flat and thin bases to enable capacitive coupling between the touchscreen electrodes and the liquid sample. However, we note that containers with other base geometries—such as curved or irregular bottoms—can produce nonuniform or distorted response patterns on the screen, while thicker bases tend to attenuate the signal and may reduce it below the touchscreen’s detection threshold.
Broader device integration. A limitation of this work is that the system has only been evaluated on a single device. This is primarily because manufacturers typically do not provide public access to raw capacitance heatmap data, which is typically available only through research collaborations. However, our work represents a first step towards what could be made possible through this sensing modality.
@@
Experimental control factors. We note that different physical factors affect the screen capacitance and need to be controlled to ensure a fair comparison between liquid samples. First, liquid capacitance varies based on temperature and humidity, Second, sample volume affects the heatmap signature of the drop given that larger volumes spread over a larger area. Thus standardized environmental conditions and operating procedures would be needed for accurate and repeatable results, however this requirements are no different than for lab-grade devices.
@@ 9.2. Future directions
Vibration motor. In initial testing, we observed that when the tablet is balanced on top of a surface such as the motorized tilt platform in Fig. 23, the vibration motor generates sufficient force to agitate and gradually move small objects such as a 10 g weight (wei, 2025), across the screen. This occurs when the material and geometric properties of the object resonates with the motor’s fundamental frequency. We note that the object does not move when the tablet is placed flat on the table, which dampens the effect of the vibration.

### v4

@@ 7.2. Concentration detection
We find that in the case of DNA and salt water (Fig. 13a,c) our system is primarily effective as a binary detector of trace concentrations of liquid. We note that in these experiments the touchscreen measures the combined capacitance of both the analyte and the buffer solution. However, any changes in capacitance across dilutions are due to varying analyte concentration. Specifically in the case of DNA it is able to consider concentration of 0–20 ng/ as a low concentration and 40 ng/ as a high concentration and distinguish between those, essentially functioning as a detector with a threshold of 20 ng/, yielding an overall detection accuracy of 95.6%. Similarly, in the case of salt water, the touchscreen is able to function a detector of the liquid with a threshold of M with an overall detection accuracy of 95.7%. In the case of ethanol, we find that our system is able to distinguish between six different concentrations of alcohol specifically 0, 10, 30, 50, 80, 100% with an overall accuracy of 94.8% (Fig. 13b).
@@ 7.4. Benchmarks
Cross-day validation. We conduct a cross-day evaluation using the multi-class ethanol concentration classification task described in Sec. 7.2. In this evaluation, training and testing data are collected on different days separated by a 74-day interval, thereby capturing both session-to-session variation and longer-term environmental changes. All test samples originate from an unseen measurement day and unseen sessions. In both training and testing, droplet samples are placed at randomized locations across the screen, to capture variations in sensitivity across spatial regions. The dataset composition for training and testing is summarized in Table A.4.
As shown in Fig. 16, the cross-day evaluation achieves an overall classification accuracy of 93.1%, comparable to the same-day training setting of 93.0% shown in Sec. 7.2. These results suggest that the system maintains stable performance over extended time intervals and is robust to session-level variability and day-level environmental noise.
| Device | Pitch size (mm) | Controller | Touch integration technology |
| Samsung Galaxy Note 10.1 | 4.2 | Atmel mXT1664s | On-cell |
| LG Nexus 5 | 4.1 | Synaptics S3350B | In-cell |
| ATEVK-MXT2952TD-A maXTouch Evaluation Kit | 4.8 | Microchip ATMXT2952TD | N/A |
Cross-device characterization. To characterize our system’s behavior across different devices, we tested three devices with different specifications: the Samsung Galaxy Note 10.1 tablet, the LG Nexus 5 smartphone, and the ATEVK-MXT2952TD-A maXTouch Evaluation Kit. While all three devices have similar pitch sizes (4–5mm), they use different touch controllers and different touch integration technology (Table 2).
We deposited liquid drops of various volumes on each device and recorded the raw capacitance data. We show in Fig. 17 the capacitance response of different devices for different volume levels. The figure shows that each device reports different absolute capacitance values, which is expected since touch controllers use arbitrary device units that are not standardized. However, deionized water and tap water, which differ significantly in conductivity, remain separable across all devices, and the trends are consistent. Due to the Nexus 5’s lower sensitivity, it cannot reliably detect drops below 50 µL for tap water and 100 µL for deionized water (Fig. 17a). This suggests that these different touchscreens are all able to perform liquid sensing albeit to different degrees.
We observe that the touch integration technology used by each device affects its sensitivity for liquid sensing. In-cell displays integrate touch sensing within the LCD layer, requiring the capacitive signal to pass through the liquid crystal material, which reduces sensitivity. On-cell displays place the touch sensor above the LCD but below the cover glass, providing better sensitivity. The external evaluation kit has no display stack, offering direct capacitive coupling and the highest sensitivity. The difference in sensitivity is also reflected in Fig. 18: for identical 200 µL drops of deionized and tap water, the external screen shows the largest capacitance change, followed by the Samsung tablet (on-cell), and then the Nexus 5 (in-cell).
Finally, the standard priming drop technique works effectively on both the Samsung tablet and external screen. The Nexus 5’s lower sensitivity requires a modified approach: spreading a small water drop with a finger to form a thin conductive film. While the execution differs, the underlying physics principle is the same: creating a conductive layer that mimics a persistent touch input.
@@ 8.1. Limitations
Container material, geometry, and thickness. The containers that we selected for our through-container sensing experiments were made of plastic and had flat and thin bases to enable capacitive coupling between the touchscreen electrodes and the liquid sample. However, we note that containers made of other materials (e.g. glass, ceramic, metal) may exhibit different dielectric properties and require material-specific calibration. Furthermore, containers with irregular base geometries can produce distorted response patterns on the screen, while thicker bases may reduce the signal below the touchscreen’s detection threshold.
Broader device integration. While our main system has been developed and evaluated primarily on one device, we characterized the capacitive response across three devices with different controllers and different touch integration technologies. Our cross-device experiments demonstrate that the underlying physics principles remain consistent across platforms, despite variations in absolute capacitance values and sensitivity. This suggests that while device-specific calibration is necessary, the sensing approach can be adapted to different touchscreen devices.
@@
Sample drop volume. Our experiments used a sample drop volume of 500 µL because it produces spatial patterns large enough relative to the touchscreen’s electrode grid to make different liquids’ capacitance features distinguishable. While the smallest detectable sample volume in our setup is approximately 10 µL, feature separability decreases as sample volume drops. Using touchscreens with denser electrode grids may improve separability at smaller volumes.
Horizontal, flat surface as operating assumption. Our system requires that the tablet be placed on a horizontal, flat surface, which is a reasonable assumption for our envisioned application scenarios in mostly controlled environments. This is because generalizing across tilting angles is an ill-posed physics problem. Specifically, even seemingly identical liquid drops on a uniform surface will exhibit different deformation and motion patterns when tilted due to contact-angle hysteresis and microscopic surface variations (Eral et al., 2013; Bonn et al., 2009). As a result, it is challenging to practically compensate for tilt-induced behavior, which motivates our explicit operating assumption of a level surface.
Cross-temperature compensation. Temperature compensation must be liquid- and concentration-specific. Our system’s response depends on three properties: conductivity, permittivity, and surface tension. In our benchmark experiment (Sec. A.4) we are able to use a linear model to compensate between tap and deionized water as both these liquids differ primarily in conductivity only, and conductivity is known to produce linear changes in temperature (Hayashi, 2004). However, for liquids which differ in permittivity and surface tension even at ambient temperature, the situation is more complex as the capacitance readings will vary nonlinearly with temperature (Uematsu and Franck, 1980; Persson, 2017; Vázquez et al., 1995).
Screen protectors. As of 2024, more than 68% of active smartphones globally use screen protectors (scr, 2026). We tested a commercially available screen protector (approximately 120 µm thick) and found that it attenuates the screen’s capacitive response to liquids (Figure 9(b)). The signal attenuation is significant enough that the priming drop does not function reliably. However, users can still perform measurements by maintaining continuous finger contact with the screen throughout the measurement process, which provides the necessary capacitive coupling. While this workaround enables functionality, it does reduce the convenience of our approach for users with thick screen protectors.
Priming. The priming mechanism does not specifically require deposition tools such as a pipette or eyedropper. Users can simply wet the screen with a fingertip and remove the priming drop using a tissue via capillary action (Fig. 4). Furthermore, priming is not strictly necessary, as users can maintain continuous finger contact on the screen to suppress the adaptive filter. While custom touchscreens and controllers are a viable alternative, our contribution lies in demonstrating liquid sensing on unmodified touchscreens.
@@ 8.2. Future directions
Vibration motor. In initial testing, we observed that when the tablet is balanced on top of a surface such as the motorized tilt platform in Fig. A.8, the vibration motor generates sufficient force to agitate and gradually move small objects such as a 10 g weight (wei, 2025), across the screen. This occurs when the material and geometric properties of the object resonates with the motor’s fundamental frequency. We note that the object does not move when the tablet is placed flat on the table, which dampens the effect of the vibration.
@@
Appendix A Appendix
A.1. Primer on capacitive sensing on touchscreens
Modern multi-point touchscreens (Fig. A.1a) are comprised of multiple layers starting from bottom to top:
Touchscreens operate on the idea of mutual capacitance which forms between the two electrode lines (Fig. A.1b) (Barrett and Omote, 2010; Hessar et al., 2016). The screen structure is composed of a 2-D array of electrodes from the driving and sensing lines. On the driving line is an AC input drive signal , which creates a shared electric field between the layers, this is sensed by the other layer with the sensing electrodes. We show in Fig. A.1c the driving signal of Samsung Galaxy Note 10.1 captured on an oscilloscope (T3DSO1302A, Teledyne LeCroy). The signal is captured by placing the oscilloscope probe directly in contact with the top of the touchscreen and connecting the ground to a metal pole. The driving frequency is approximately 100 kHz with a peak to peak distance is approximately 2.4 mV.
In a touchscreen without any object in contact with it, a mutual capacitive field forms between the driving and sensing electrodes (Fig. A.2a). When a conductive object is near, it disturbs the shared electric field. In the case of grounded conductive objects like a human finger that is coupled to Earth ground, the electric field lines are pulled towards the finger and it reduces the capacitance between two electrodes at an intersection (Fig. A.2b). The touchscreen controller is able to read the change in capacitance and detect where it has occurred. The human body serves as an effective ground because it is capacitively coupled to Earth, forming a large conductive body with a low-impedance pathway to ground. For the electric field, this provides an easier route to discharge, as the finger is in closer proximity to the screen compared to some of the sensing electrodes which can be farther away and have higher impedance.
In contrast, when placing a liquid sample on the screen it alters the electric field coupling between the driving and sensing electrodes in two ways (Fig. A.2c):
(A.1) where is electrical conductivity of the liquid, is the electrode area, is the separation distance or pitch between electrodes, and is the angular frequency of excitation . The remaining charge is redistributed to neighboring electrodes, effectively reducing the mutual capacitance readings (Zimmerman et al., 1995).
(A.2) where is vacuum permittivity. Non-conductive or poorly conductive liquids such as oils and alcohols influence the measured capacitance primarily through this effect creating an increase in the measured capacitance.
A.2. Empirical characterization and analysis
A.2.1. Multi-scale spatial sensitivity analysis
To characterize the change in capacitance instead of using the centroid which only measures the value at a single cell, which would make sense if the droplet was deposited only in the center. For drops in other locations which straddle the boundary between multiple cells we sum together the values across cells with a negative change in capacitance. We show in Fig. A.3a the magnitude capacitance of the droplet. Both the 10 and 25 samples had comparatively higher capacitance values and lower variation at the center and vertical locations, and lower values at the corner and horizontal locations. The results suggest that placing droplets at the center of the electrode produce the most reliable results. This is the position that we adopt in all experiments. We note however that the use of larger volume samples occupy multiple cells and our system captures features of the spatial information in the region around the centroid of the drop for downstream applications.
with , , and .
The sensitivity map in Fig. A.3b, shows two clusters to the left and right of the screen which are sensitive, while the sides and the center have low sensitivity.
A.2.2. Temporal noise characterization and reduction
Here, we characterize the touchscreen’s sensor noise over time. We note that because the touchscreen is primarily designed to detect touch locations rather than precise capacitance values, it is inherently not optimized for low-noise measurements. To characterize the noise profile, we record the tablet capacitance readings over the course of 30 seconds without anything on the screen. We show in Fig. A.4a the standard deviation per pixel across the full screen. We can observe the standard deviation is high in the center of the screen, and lower at the sides, the average standard deviation is 8. To reduce the effect of noise we perform averaging across each frame. Fig. A.4b shows the effect of averaging across 20 frames (12 seconds), with an average standard deviation across the screen of 2.8. The relationship between averaging time and the average standard deviation across the screen is shown in Fig. A.4c, the plot shows the theoretical relationship of , where is the number of frames being averaged over as well as the measured decrease in standard deviation over 30 seconds. The plot shows that the measured readings decrease with a trend closely following the theoretical model.
A.3. Data collection and protocol.
| Count | Unspiked | Spiked | Total |
| # drops | 77 | 61 | 138 |
| # sessions | 7 | 6 | 13 |
| Count | Unadulterated | Adulterated | Total |
| # drops | 88 | 72 | 160 |
| # sessions | 8 | 7 | 15 |
| Count | Unadulterated | Salt-adulterated | Detergent-adulterated | Total |
| # drops | 44 | 33 | 41 | 118 |
| # sessions | 4 | 3 | 4 | 11 |
| Count | 0–20 ng/L | 40–80 ng/L | Total |
| # drops | 124 | 165 | 289 |
| # sessions | 12 | 16 | 28 |
| Count | 0% | 10% | 30% | 50% | 80% | 100% | Total |
| # drops | 77 | 122 | 78 | 159 | 87 | 9 | 532 |
| # sessions | 7 | 12 | 8 | 16 | 8 | 1 | 52 |
| Count | – M | – M | Total |
| # drops | 110 | 121 | 231 |
| # sessions | 10 | 12 | 22 |
| Count | Deionized water | Ethanol | Tap water | Total |
| # samples | 20 | 16 | 18 | 54 |
| # sessions | 5 | 4 | 5 | 14 |
| Count | Unspiked | Spiked | Total |
| # samples | 16 | 16 | 32 |
| # sessions | 4 | 4 | 8 |
| Count | 0.001% | 0.01% | Total |
| # samples | 20 | 20 | 40 |
| # sessions | 5 | 5 | 10 |
| Split | Count | 0% | 10% | 30% | 50% | 80% | 100% | Total |
| Training (Day 1) | # drops | 77 | 61 | 39 | 79 | 40 | 9 | 305 |
| # sessions | 7 | 6 | 4 | 8 | 4 | 1 | 30 | |
| Testing (Day 74) | # drops | 44 | 36 | 36 | 39 | 44 | 43 | 245 |
| # sessions | 4 | 4 | 4 | 4 | 4 | 4 | 24 |
We summarize here the data collection protocol used across all tasks in the paper, including the total number of unique liquid samples (drops), the number of data collection sessions and days, the cleaning procedures used to prevent cross-contamination, and the specifications of containers used for through-container sensing experiments. All datasets consist of physically distinct liquid drops; no synthetic data augmentation is used.
A.3.1. Data collection sessions and spatial placement
Data was collected across multiple sessions. A session is defined as a continuous data collection period from placing liquid drops or containers on the touchscreen to their removal. In all experiments, liquid drops and containers were placed randomly and uniformly across the touchscreen surface to avoid bias toward specific screen locations and to expose the model to spatial non-uniformities of the touchscreen.
The number of samples and sessions for each task is summarized in Tables A.1– A.3. Table A.1 summarizes the data composition for each adulteration detection task. Table A.2 summarizes the data composition for each adulteration detection task. Table A.3 summarizes the data composition for each adulteration detection task.
To evaluate robustness over extended time spans, we conduct a cross-day evaluation using the ethanol concentration classification task. The model is trained exclusively on data collected on one day and evaluated on data collected 74 days later, with no overlap in sessions or samples between training and testing. All samples in both training and testing are placed randomly across the touchscreen surface. Dataset composition for this evaluation is summarized in Table A.4.
A.3.2. Cleaning and Cross-Contamination Prevention
To prevent cross-contamination between samples, the touchscreen surface and any containers were cleaned between consecutive measurement sessions. After each session, the screen surface was wiped with paper towel wipes and isopropyl wipes, and allowed to dry completely before the next measurement. Containers used in through-container experiments were rinsed with deionized water and dried between samples.
A.4. Benchmarks
Measurement stability over time. We evaluate the stability of capacitance readings for the priming drop and the test liquid samples on the tablet screen. We show in Fig. A.5a that the measured capacitance at the centroid of the priming drop remains stable across the measured period of ten minutes with a mean and standard deviation of 376 7 device units. We also perform this measurement for water samples of different volumes across an 10 minute period and show in Fig. A.5b that these values remain stable across the measured period with standard deviations ranging from 6 to 10 device units across the measurement period. There is no substantial relationship between the standard deviation of the measurements with the volume of the liquid drops. We note that stability across these time scales is sufficient for averaging the effects of noise and for recording a measurement.
| Round-shaped cup | Heart-shaped cup | |
| Container dimensions | ![]() |
| Base thickness (mm) | 0.6 | 1.1 |
Cross-container generalization. To evaluate the robustness of our system across different container types, we conducted a benchmark study using two types of plastic containers with base geometries and base thicknesses.
We selected two containers (Table A.5): (1) Round plastic cup: cylindrical geometry, 0.6 mm base thickness; (2) Heart-shaped cup: non-circular geometry, 1.1 mm bottom thickness. Both containers feature thin, slightly concave bases that rest on the touchscreen surface with the outer rim making full contact.
As containers with different shaped bases have varying dimensions, direct comparison in the image space is challenging. To address this, we extract a single capacitance feature from the image of the container (mean of all positive pixels) to represent a sample. We show in Fig. A.6a that this feature is sufficiently predictive at distinguishing between different liquids.
The figure also shows that the base thickness has a direct effect on the attenuation of the capacitive signal. The heart-shaped container, with nearly twice the bottom thickness (1.1 mm vs 0.6 mm), exhibits lower capacitance values across all liquids. Importantly, the relative ordering of liquid classes remains consistent across containers. These results suggest that compensation strategy could enable cross-container generalization.
To compensate for thickness-induced signal attenuation, we derive a label-preserving transformation. Let and denote feature vectors from the training and testing containers, respectively, with thicknesses and .
The compensation ratio is defined as:
Rather than uniformly scaling all features by (which would artificially increase variance), we apply an additive offset that adjusts only the distribution’s center:
where is the mean feature vector computed over all test samples. This formulation ensures that the compensated mean aligns with the training distribution, while the variance remains unchanged, maintaining class separability.
We trained a Random Forest classifier on one container type and evaluated it on the other. Without compensation, the model fails to generalize (33.3% accuracy equals random chance). With thickness-based compensation, cross-container accuracy reaches 86.5 to 88.3%, demonstrating effective domain adaptation. Fig. A.6b and Fig. A.6c shows the confusion matrix for the compensated cross-container classification.
These results demonstrate that our system can generalize to unseen container bases with different thicknesses and shapes. The shape of the container can be controlled for by distilling the capacitance image into a single scalar feature, while container thickness introduces a systematic bias in absolute signal magnitude, which can be corrected through simple physics-based compensation.
Effect of temperature. We examine the effect of temperature on measured capacitance using tap water and deionized water. We heated up the liquid samples in a tabletop water bath (ivy, 2025) from 20–90∘C (10∘C intervals). In Fig. A.7b, the solid lines for tap water (blue) and deionized water (red) show that the measured capacitance increases for both liquids, which is due to conductivity increase as temperature increases.
Given the observed temperature effect, we investigate whether a compensation algorithm can maintain classification accuracy across temperatures. For each liquid, we split the data in half: the first half fits a linear model (where is temperature and is measured centroid capacitance), and the second half is used for evaluation.
The compensation offset at temperature is , normalizing all measurements to 20∘C. Fig. A.7b shows the compensation effectively flattens the temperature-dependent drift for both liquids. To evaluate cross-temperature generalization, we trained a logistic regression classifier on centroid features at each temperature and tested on all other temperatures. The compensation achieves 86.1% overall accuracy (69.7% without compensation).
Effect of device tilt. We conducted a benchmark study to evaluate the effect of device tilt on our system. We placed the tablet on a motorized tilt platform (til, 2025) (Fig. A.8a). We deposited ten drops of water and isopropyl alcohol (IPA) onto the surface and varied the tilt angle at increments of approximately 1 to 3∘. We show in Fig. A.8b how the droplets displace from their original position as tilt angle increases. In the case of water, which has a surface tension of , the droplets start moving at 5∘ and this displacement increases till its roll-off angle of 11.5∘ when it slides off the tablet. In contrast, IPA, which has a lower surface tension of , has a much higher rate of displacement with a roll-off angle of 2∘. While water had a displacement of approximately 4 cells at the roll-off angle, alcohol had a displacement around three times higher at 12 cells. The behavior of the drops at the rest and roll-off angles are visualized in Fig. A.8c.
Physics-informed model. We perform a benchmark study using physics-informed features that capture different aspects of the liquid’s capacitive response and Random Forest classification. We show in Table A.6 the list of selected features and their physical interpretation.
We evaluated the physics-informed model on three adulteration detection tasks. Fig. A.9 shows the confusion matrices. The model achieves 84.5% accuracy on soda spiking detection, 81.7% on wine adulteration detection, and 95.6% on milk adulteration detection. These results demonstrate that physics-informed features alone can achieve reasonable classification performance.
| Feature | Physical Interpretation | Feature importance (%) |
| Spatial spread | Weighted variance of sensor positions; measures the effective contact area influenced by surface tension | 25.2 |
| Skewness | Asymmetry of capacitance distribution; indicates non-uniform drop spreading or edge effects from wetting properties | 11.7 |
| Minimum value | Peak capacitance response; indicates the strongest electrical coupling point, influenced by liquid conductivity, permittivity and drop geometry | 9.6 |
| Centroid value | Capacitance at the weighted center of the drop; directly reflects permittivity and conductivity at the point of maximum coupling | 8.8 |
| Standard deviation | Variability in capacitance across the drop; relates to non-uniformity in drop shape and thickness from surface tension effects | 8.4 |
| Mean absolute value | Average magnitude of capacitance response; proportional to overall liquid permittivity and conductivity | 8.3 |
| 25th percentile | Lower quartile capacitance; characterizes the distribution of strong coupling regions | 8.3 |
| Value range | Difference between maximum and minimum; related to the conductivity and permittivity of the liquid | 7.5 |
| 75th percentile | Upper quartile capacitance; characterizes the distribution of weak coupling regions | 7.3 |
| Maximum value | Peak capacitance response; captures the fringing effect of the liquid drop | 5.1 |
However, the accuracies are 5-10 percentage points lower than the CNN-based approaches, suggesting that while physics-informed features provide a strong foundation, the CNN can learn additional subtle patterns from the raw sensor data that are difficult to manually specify. That being said, further feature engineering could be performed to develop a hybrid model combining physics-informed features, and data-driven features for use by a downstream classifier.
Feature importance analysis. An important advantage of using physics-informed features is its interpretability. We perform a feature importance analysis which reveals that spatial spread (25.2%) and skewness (11.7%) are the most discriminative features, both of which primarily correspond to surface tension effects that affect drop spreading and wetting behavior. Following these, the minimum value (9.6%) and centroid value (8.8%) have the most discriminative value, and they are associated with the liquid’s conductivity and permittivity.
Ablation study. We perform an ablation study on the liquid adulteration datasets to examine the effect of different system design choices on performance (Fig. A.10). Specifically, we examine the effect of number of training samples, region size around the centroid, and machine learning model on system performance. For these experiments, we consider a subset of the adulteration classes: for soda spiking and wine adulteration we compared to ethanol of 10, 20, 30% as a matter of computational efficiency to examine the overall effect of the system parameters. Here we used 500 drop sizes.
For the training dataset size (Fig. A.10a) we evaluated the affect of changing the size of the training set from 0, 25, 50, 75, 100% of the original size. We find that as expected this increases the accuracies with diminishing returns. In the soda and wine dataset, we find that there are no substantial performance increases beyond approximately 200 training samples.
For the region size (Fig. A.10b), we find that accuracy increases steadily as the region size increases from a patch to a patch. In the case we use a fully-connected network instead of a CNN. For larger classes we use an adaptive CNN where for and classes there are only two convolutional layers without any pooling, while for larger patch sizes of to there is a single max pooling layer, and for there are 2 pooling layers. This result is expected as the screen area covered by a drop can expand to affect the electric field of a grid, and a larger grid captures this information.
Finally, we examined the effect of model architecture on system performance (Fig. A.10c). We find that using a CNN only, LSTM only, CNN+LSTM and CNN+Transformer provide performance within a similar range of 91 to 94%. Tree-based models such as XGBoost and Random Forest provide lowered accuracies of 79 to 81%. The results suggest that global features about the region, can be captured by the tree-based models. However, deep learning modules better capture the non-linear relationship (Sec. 4) between classes, with the CNN and LSTM producing higher accuracies. As CNN alone and LSTM alone are able to produce high accuracies, it suggests that non-linear feature combinations are sufficient to perform classification.
Temporal length sensitivity analysis. We evaluate the effect of temporal context (number of training frames) on system performance for different liquid sensing tasks. We trained classifiers using sequences of 1, 2, 5, 10, 20, and 50 frames, with all frames provided as independent training samples. At test time, each frame is classified independently. As each frame corresponds to a 0.6-second interval, our deployed system achieves an inference latency of 0.6 seconds.
Fig. A.11 shows classification accuracy as a function of training temporal window length across all tasks. While we used 50-frame sequences (30 seconds) in our primary experiments to maximize accuracy, the results demonstrate that substantially shorter training sequences suffice for most applications.
Specifically, adulteration detection tasks (coke spiking, milk adulteration) achieve ¿95% accuracy when trained on short sequences of 5 frames (3 seconds), and concentration detection of ethanol maintains ¿90% accuracy at this temporal resolution. The through-container task, which involves sensing through opaque plastic barriers, requires longer training sequences, reaching 90% accuracy when trained on 20-frame sequences (12 seconds).
Our design classifies each frame independently instead of averaging measurements over time windows, which provides three key advantages:
Low inference latency: Our models can classify liquids in real-time with 0.67-second latency per frame, enabling immediate feedback.
Data augmentation: Training on individual frames rather than averaged sequences substantially increases the training dataset size.
Noise robustness: By training on noisy individual frames rather than averages, our model learns to establish decision boundaries that are robust to measurement noise. This produces classifiers that can perform inference on single, noisy measurements.